OpenStudy (anonymous):

Rewrite the rational exponent as a radical by extending the properties of integers exponents. 2^3/4 / 2^1/2 the choices are 8 radical2^3 Radical 2^3/4 4 radical 2 radical2

OpenStudy (anonymous):

This is the problem right?\[\frac{ 8 }{ \frac{ 4 }{ \sqrt{2} } }\]

OpenStudy (anonymous):

No the problem is 2^3/4 divided by 2^1/2 and the choices are the rest I wrote up there

OpenStudy (jdoe0001):

\(\Large { \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}}\implies a^{-\frac{{\color{blue} n}}{{\color{red} m}}} \\\quad \\ \\ \quad \\ \cfrac{2^{\frac{3}{4}}}{2^{\frac{1}{2}}}\implies \cfrac{2^{\frac{3}{4}}}{1}\cdot \cfrac{1}{2^{\frac{1}{2}}}\implies \cfrac{2^{\frac{3}{4}}}{1}\cdot2^{{\color{red}{ -}}\frac{1}{2}}\implies 2^{\frac{3}{4}}\cdot2^{{\color{red}{ -}}\frac{1}{2}} }\)

OpenStudy (anonymous):

Here is the paperwork. Help me please!!!, can you guys tell me if the other ones are right??

OpenStudy (jdoe0001):

well... do you know how to add/subtract fractions?

OpenStudy (anonymous):

Finding a common denominator and just subtract the numerator

OpenStudy (anonymous):

Got it 4 radical 2, option 3 from up there